A magazine provided results from a poll of 2000 adults who were asked to identify their favorite pie. Among the 2000 respondents, 1 1% chose chocolate pie, and the margin of error was given as ± 5 percentage points. What values do p, q, n, E, and p represent? If the confidence level is 95%, what is the value of α? The value of p is The value of q is The value of n is the population proportion. found from evaluating 1-p the sample size. the sample proportion. the margin of error The value of E is The value of p is If the confidence level is 95%, what is the value of α? α-O (Type an integer or a decimal. Do not round.)
Answer #1

Solution :

The value of \hat p is the sample proportion = 11% = 0.11

The value of \hat q is found from the evaluating 1 - \hat p = 1 - 0.11 = 0.89

The value of the n is the sample size = 2000

The value of the E is the margin of error = 5% = 0.05

The vale of the p is the population proportion .

\alpha = 1 - 95% = 1 - 0.95 = 0.05